# JWST Insights into Narrow-line Little Red Dots

### v1

@@ Abstract
James Webb Space Telescope (JWST) has revealed a population of red and compact objects with a unique V-shape SED at known as “Little Red Dots” (LRDs). Most of the LRDs with existing spectral observations exhibit broad Balmer lines and are thus likely host active galactic nuclei (AGNs). Here we present a study of LRDs with narrow emission lines and without broad H lines. Our sample consists of five LRDs at with H line widths of about 250 (except one around 475 ). They are selected from 32 LRDs that have NIRSpec high- or medium-resolution grating spectra covering H. During our construction of the sample, we find that of the LRD candidates previously selected do not show red continua emission but resemble the V-shape spectra due to strong line emission. We explore various properties of the five narrow-line LRDs using JWST spectral and imaging data. Compared to normal star-forming galaxies, they tend to have relatively higher H line widths and luminosity but smaller equivalent width. If these LRDs are dominated by galaxies, our SED modeling suggests that they are dusty, compact star-forming galaxies with high stellar masses and star formation rates (SFRs). Alternatively, if their SEDs are produced by AGNs, the inferred central black hole masses () are in the range of –, placing them at the low-mass end of the AGN population. They may represent an early stage of super-Eddington growth, where the black holes have yet to accumulate significant masses. With large measurement uncertainties, these black holes appear slightly overmassive relative to the local – relation, but consistent or undermassive with respect to the – and – relations. We further find that nearly half of the high-redshift broad-line AGNs exhibit V-shape SEDs. Our narrow-line LRDs represent a distinct population and offer new insights into the diversity of LRDs.
1 Introduction
@@
One of the most intriguing results from early observations with the James Webb Space Telescope (JWST) is the identification of an abundance of red and compact objects (e.g., I. Labbé et al., 2023; G. Barro et al., 2024). These objects exhibit a unique “V-shape” spectral energy distribution (SED) that is red in the rest-frame optical and blue in the rest-frame ultraviolet (UV). They are called “Little Red Dots” (LRDs) for their compact morphology (typical pc; e.g., J. F. W. Baggen et al., 2023; L. J. Furtak et al., 2023; C. A. Guia et al., 2024) and red colors in the observed-frame – range (e.g., J. Matthee et al., 2024). LRDs are ubiquitous at redshifts from up to (e.g., G. Barro et al., 2024; G. C. K. Leung et al., 2023; D. D. Kocevski et al., 2024; V. Kokorev et al., 2024a), while at lower redshift their number density decline rapidly (e.g., D. D. Kocevski et al., 2024; Y. Ma et al., 2025). Spectroscopy of more than one hundred photometrically identified LRDs reveals that of them exhibit broad Balmer lines (e.g., Y. Harikane et al., 2023; J. E. Greene et al., 2024; M. Killi et al., 2024; D. D. Kocevski et al., 2024; J. Matthee et al., 2024; B. Wang et al., 2024).
@@
The physical origins of the V-shape SED and the nature of LRDs remain unclear. A natural explanation of the broad lines and compact sizes is that LRDs are active galactic nuclei (AGNs). However, the presence of prominent Balmer breaks observed in several LRDs (e.g., J. F. W. Baggen et al., 2024; V. Kokorev et al., 2024b; B. Wang et al., 2024), along with the general lack of X-ray (e.g., T. T. Ananna et al., 2024; M. Yue et al., 2024; R. Maiolino et al., 2025), mid-IR emission (e.g., P. G. Pérez-González et al., 2024; C. C. Williams et al., 2024), and variability signatures (e.g., M. Kokubo & Y. Harikane, 2024; W. L. Tee et al., 2025; Z. Zhang et al., 2025), pose challenges to the AGN interpretation. These issues, however, may be mitigated under a scenario involving super-Eddington accreting black holes enshrouded by extremely dense gas (e.g., K. Inayoshi et al., 2024; K. Inayoshi & R. Maiolino, 2025; X. Ji et al., 2025; R. P. Naidu et al., 2025), supported by the high fraction () of Balmer absorption in high- and medium-resolution spectra (e.g., X. Lin et al., 2024; A. J. Taylor et al., 2024; I. Juodžbalis et al., 2025). In this context, LRDs may represent the very first activity of supermassive black hole (SMBH) growth (K. Inayoshi, 2025), potentially linked to the formation of seed black holes via non-standard seeding channels, such as the core-collapse of self-interacting dark matter halos (F. Jiang et al., 2025).
@@
While the majority of spectroscopically confirmed LRDs exhibit broad emission lines, it remains unclear whether the characteristic V-shape continuum is necessarily associated with broad-line emission. Some broad emission lines of LRDs were identified using low-resolution () NIRSpec/PRISM spectra (e.g., J. E. Greene et al., 2024), and their intrinsic line widths can be much smaller. For instance, one LRD with a broad H FWHM reported by J. E. Greene et al. (2024) was later found to be well fitted by a single Gaussian with FWHM in higher-resolution () NIRCam/WFSS slitless spectra (S. Fu et al., 2025). More recently, J. Zhang et al. (2025) report that 68% of color-selected LRDs with H detections in the NIRCam/Grism spectra in the GOODS-N field do not exhibit broad-line features. Since the detection of broad components in these slitless spectra is limited by sensitivity, deeper and higher-resolution NIRSpec data are needed to confirm their absence. Nevertheless, these results suggest that a subset of LRDs may exhibit only narrow lines while sharing key photometric properties with their broad-line counterparts, such as compact sizes and V-shape SEDs. Understanding the physical origins of narrow-line LRDs may provide crucial insights into the formation pathways and duty cycles of LRDs.
@@
The situation is further complicated by the lack of a clear definition and robust selection methodology for LRDs in the literature. A commonly adopted definition identifies LRDs as sources with compact morphologies at long wavelengths and V-shape SEDs. Some studies further require the presence of broad emission lines as part of the definition. In practice, many sources are classified as LRDs based on simple color-cut or slope-cut criteria (e.g., H. B. Akins et al., 2024; D. D. Kocevski et al., 2024; V. Kokorev et al., 2024a; I. Labbe et al., 2025). These photometric colors and fitted slopes may be contaminated by strong emission lines, such as H, H, and [O iii] line (e.g., K. N. Hainline et al., 2025). This leads to misclassifications or the inclusion of unrelated sources in LRD samples. Therefore, it is crucial to remove the contamination of the line emission when constructing a LRD sample.
@@
In this work, we present a systematic investigation of narrow-line LRDs based on a clean sample selected using emission-line-free photometry. The contamination from strong emission lines is removed using spectroscopic information from NIRSpec/PRISM observations. The narrow H line is identified with high- or medium-resolution NIRSpec grating spectra. We compare the properties of narrow-line LRDs with those of broad-line LRDs and normal star-forming galaxies, aiming to understand the diversity within the LRD population and the possible links between them. The paper is organized as follows. In Section 2, we describe our data and data reduction. In Section 3, we introduce our selection and characterization of the narrow-line LRD sample. Two possible interpretations of the narrow-line LRDs are proposed in Section 4. In Section 5, we discuss the comparison of the sample selection methods and the implication of narrow-line LRDs. Conclusions are given in Section 6. Throughout this paper, we use a cosmology with , , and (Planck Collaboration et al., 2020). By default, measurement uncertainties are quoted at a 1 confidence level, and upper limits are quoted at a 90% confidence level.
2 Data Description and Reduction
2.1 Imaging Data and Catalog
@@
We collect publicly available JWST/NIRCam data in A2744, CEERS, GOODS-S, GOODS-N, COSMOS, and UDS fields. The detailed reduction steps can be found in Z. Zhang et al. (2025). Briefly, we reduce the raw data using the combination of the JWST Calibration Pipeline (v.1.12.5), the CEERS NIRCam imaging reduction (M. B. Bagley et al., 2023),111https://github.com/ceers/ceers-nircam. and our own custom codes. All NIRCam images are aligned to the HST images in each field. We use the public HST image from the CANDELS and 3D-HST surveys (N. A. Grogin et al., 2011; A. M. Koekemoer et al., 2011; G. B. Brammer et al., 2012; I. G. Momcheva et al., 2016). The final mosaics in all fields have pixel scales of .
@@
The empirical PSF models of each field are constructed from the mosaic images with PSFEx (E. Bertin, 2011), following the method in M.-Y. Zhuang & Y. Shen (2024). We then use these PSFs to derive kernels to match all PSFs to the F444W PSF using PyPHER (A. Boucaud et al., 2016). These kernels are used to convolve all images to match the F444W image. Then we perform photometry using SExtractor (E. Bertin & S. Arnouts, 1996) on the PSF-matched images to construct source catalog. We use the dual image mode and the detection image is an image combining the PSF-matched F277W, F356W and F444W images. We adopt a two-step ‘cold + hot’ mode to achieve an optimal source detection (M. Barden et al., 2012). Photometry is measured in a small Kron aperture (=1.1, =1.6), and an aperture correction is performed to match the default Kron aperture (=2.5, =3.5) in each band. Following the method in M. J. Rieke et al. (2023), we measure the photometric errors by placing 100,000 random apertures in the source-free region of the images. These apertures are divided into 100 groups based on their weight values. For each group, a power-law relation between aperture size and flux scatter is fitted and used to derive the sky background noise for each source according to its weight value and aperture size.
@@
For the CEERS field, we adopt the ASTRODEEP catalog (E. Merlin et al., 2024) that is based on the images from the CEERS data release (M. B. Bagley et al., 2023), since this catalog includes more HST bands. Their basic procedure for photometry is similar to ours, with the major differences outlined below. Source detection was carried out using SExtractor on a weighted combination of the F356W and F444W images. The total fluxes were inferred within a preferred aperture, determined via the object’s segmentation map. The aperture correction in the detection band was applied to other bands under the assumption of no color gradient.
2.2 NIRSpec Spectroscopic Data
@@
The NIRSpec spectroscopic data used in this work are from the DAWN JWST Archive (DJA222https://dawn-cph.github.io/dja/.), which includes fully reduced 1D and 2D spectra of all public JWST/NIRSpec data. We use the version 3 publicly released NIRSpec datasets, including 26,077 spectra for 14,909 sources from 43 programs. The source IDs used in this paper are the UIDs of the DJA dataset. These data are reduced using the latest version of msaexp and the JWST Calibration Pipeline (v.1.14.0) with calibration files jwst_1225.pmap from the Calibration Reference Data System. The detailed MSA data reduction procedures using msaexp can be found in A. de Graaff et al. (2024a) and K. E. Heintz et al. (2024). The absolute flux calibration accuracy for NIRSpec MSA spectroscopy is generally at a level of 10% to 20% (A. de Graaff et al., 2024a). DJA also offers spectroscopic redshifts obtained through the template fitting algorithms employed in the msaexp pipeline. These redshifts are visually inspected and graded for reliability, with grade 3 indicating a secure measurement. In our analysis, we only include sources with grade 3 redshift, and adopt the grade 3 redshift as their redshift value.
@@
We also notice the availability of two large NIRSpec datasets from the data releases of two JWST Cycle 1 programs: JWST Advanced Deep Extragalactic Survey (JADES; D. J. Eisenstein et al., 2023; F. D’Eugenio et al., 2024) and the CEERS (M. B. Bagley et al., 2023). These datasets are carefully reduced and generally offer higher data quality compared to the DJA data (e.g., L. Jiang et al., 2025). For uniformity, we adopt only the DJA data throughout this work. We have verified that this choice has a negligible effect on our results. Most of the CEERS spectra are also included in DJA. In the JADES release, 1092 out of 2525 targets are not available in DJA; however, we identified only 4 LRDs (and no narrow-line LRDs) among them, which may be attributed to the target selection strategy of the JADES survey. Overall, the results obtained from the JADES/CEERS and DJA datasets are consistent for common targets.
3 Narrow-line LRDs
3.1 Selection of LRDs
@@
Photometric selection of LRDs in the literature primarily relies on their observed colors to identify the characteristic V-shape SEDs (e.g., H. B. Akins et al., 2024; V. Kokorev et al., 2024a; I. Labbe et al., 2025). D. D. Kocevski et al. (2024) further refined this approach by applying cuts to the rest-frame optical and UV slopes. This photometry-only selection may suffer from inaccurate photo- or boosted red colors by strong emission lines (e.g., K. N. Hainline et al., 2025).
@@
LRDs are then selected using the criteria of and , as done by D. D. Kocevski et al. (2024). We present two examples in Figure 1 illustrating sources with a fake and a genuine V-shape continuum, respectively. For comparison, we also show the LRD template constructed from the average SED of LRDs in H. B. Akins et al. (2024). In Figure 1(a), the red color of the source is primarily due to its strong H and [O iii] emission lines. After subtracting their contributions, the optical slope becomes flat. In contrast, the source shown in Figure 1(b) remains its V-shape after the emission line contribute is subtracted. The distribution of the best-fit UV and optical slopes for our sample is shown in Figure 2(a). In total, there are 179 sources with and .
To select sources with compact morphology, we apply an additional cut on the half-light radius measured by SExtractor in the F444W-band images. We derive a relation between and for point sources, as shown in Figure 2(b). Sources with below 1.5 times the relation (green dashed line in Figure 2b) are selected. A total of 98 LRDs meet this criterion. The redshift-magnitude diagrams (F115W and F444W) of the selected LRDs are shown in Figure 3. These LRDs span a large redshift range of –9 and a broad magnitude range of –28 mag. In the rest-frame UV, their magnitudes are comparable to those of typical star-forming galaxies reported by M. Curti et al. (2024), whereas in the rest-frame optical, the LRDs appear brighter.
@@
3.2 Selection of Narrow-line LRDs
We aim to search LRDs that do not have a broad H line component. Among the 98 selected LRDs, 32 have high- or medium-resolution grating spectra covering H, and 20 of them have additional high- or medium-resolution grating spectra covering H and O iii. We fit these lines using lmfit (M. Newville et al., 2025). The line spread functions (LSFs) provided by JDOX significantly underestimate the actual spectral resolution (e.g., A. de Graaff et al., 2024b; R. Maiolino et al., 2024), since they apply only for uniformly illuminated sources. Therefore, we use msafit to model the LSF, which forward models the light trajectory on the JWST detector for a given source position to measure the spectral resolution (A. de Graaff et al., 2024b). Some sources have multiple high- or medium-resolution spectra. In such cases, we use the spectrum with the highest S/N rather than combining them to avoid ambiguity due to the LSF effect. The results from different spectra of the same sources are consistent.
For O iii, we use a single Gaussian to fit either line separately. We do not apply any constraint (e.g., L. Jiang et al., 2025). For H and H, we perform two types of fitting:

### v2

@@ Abstract
JWST has revealed a population of red and compact objects with unique V-shape SEDs known as “Little Red Dots” (LRDs). Many LRDs exhibit broad Balmer lines and thus likely host AGNs. Here we present a study of five LRDs without broad H lines at . They are selected from 32 LRDs that have NIRSpec high- or medium-resolution grating spectra covering H, and their H line widths are around 250 . Compared to normal star-forming galaxies, narrow-line LRDs tend to have relatively higher H line widths and luminosities. If they are dominated by galaxies, our SED modeling suggests that they are dusty, compact star-forming galaxies with high stellar masses and star formation rates. Alternatively, if their SEDs are produced by AGNs, the inferred black hole masses (–) place them at the low-mass end of the AGN population. They may represent an early stage of super-Eddington growth, where the black holes have yet to accumulate significant masses. With large uncertainties, these black holes appear slightly over-massive relative to the local – relation, but consistent or under-massive with respect to the – and – relations. These narrow-line LRDs offer new insights into the diversity of LRDs. Finally, we find that of high-redshift broad-line AGNs exhibit V-shape SEDs. In addition, of the previous LRD candidates are not real LRDs, and their V-shape SEDs are caused by strong line emission.
I Introduction
@@
One of the most intriguing results from early observations with the James Webb Space Telescope (JWST) is the identification of an abundance of red and compact objects (e.g., 2023Natur.616..266L; 2024ApJ...963..128B). These objects exhibit a unique “V-shape” spectral energy distribution (SED) that is red in the rest-frame optical and blue in the rest-frame ultraviolet (UV). They are called “Little Red Dots” (LRDs) for their compact morphology (typical pc; e.g., 2023ApJ...955L..12B; 2023ApJ...952..142F; 2024RNAAS...8..207G) and red colors in the observed-frame – range (e.g., 2024ApJ...963..129M). LRDs are ubiquitous at redshifts from up to (e.g., 2024ApJ...963..128B; 2023ApJ...954L..46L; 2024arXiv240403576K; 2024ApJ...968...38K), while at lower redshift their number density decline rapidly (e.g., 2024arXiv240403576K; 2025arXiv250408032M). Spectroscopy of more than one hundred photometrically identified LRDs reveals that of them exhibit broad Balmer lines (e.g., 2023ApJ...959...39H; 2024ApJ...964...39G; 2024A&A...691A..52K; 2024arXiv240403576K; 2024ApJ...963..129M; 2024arXiv240302304W).
@@
The physical origins of the V-shape SED and the nature of LRDs remain unclear. A natural explanation of the broad lines and compact sizes is that LRDs are active galactic nuclei (AGNs). However, the presence of prominent Balmer breaks observed in several LRDs (e.g., 2024arXiv240807745B; 2024arXiv240720320K; 2024arXiv240302304W), along with the general lack of X-ray (e.g., 2024ApJ...969L..18A; 2024ApJ...974L..26Y; 2025MNRAS.538.1921M), mid-IR emission (e.g., 2024ApJ...968....4P; 2024ApJ...968...34W), and variability signatures (e.g., 2024arXiv240704777K; 2025ApJ...983L..26T; 2025ApJ...985..119Z), pose challenges to the AGN interpretation. These issues, however, may be mitigated under a scenario involving super-Eddington accreting black holes enshrouded by extremely dense gas (e.g., 2024arXiv241203653I; 2025ApJ...980L..27I; 2025arXiv250113082J; 2025arXiv250316596N), supported by the high fraction () of Balmer absorption in high- and medium-resolution spectra (e.g., 2024arXiv240717570L; 2024arXiv240906772T; 2025arXiv250403551J). In this context, LRDs may represent the very first activity of supermassive black hole (SMBH) growth (2025arXiv250305537I), potentially linked to the formation of seed black holes via non-standard seeding channels, such as the core-collapse of self-interacting dark matter halos (2025arXiv250323710J).
@@
While the majority of spectroscopically confirmed LRDs exhibit broad emission lines, it remains unclear whether the characteristic V-shape continuum is necessarily associated with broad-line emission. Some broad emission lines of LRDs were identified using low-resolution () NIRSpec/PRISM spectra (e.g., 2024ApJ...964...39G), and their intrinsic line widths can be much smaller. For instance, one LRD with a broad H FWHM reported by 2024ApJ...964...39G was later found to be well fitted by a single Gaussian with FWHM in higher-resolution () NIRCam/WFSS slitless spectra (2025arXiv250303829F). More recently, 2025arXiv250502895Z report that 68% of color-selected LRDs with H detections in the NIRCam/Grism spectra in the GOODS-N field do not exhibit broad-line features. Since the detection of broad components in these slitless spectra is limited by sensitivity, deeper and higher-resolution NIRSpec data are needed to confirm their absence. Nevertheless, these results suggest that a subset of LRDs may exhibit only narrow lines while sharing key photometric properties with their broad-line counterparts, such as compact sizes and V-shape SEDs. Understanding the physical origins of narrow-line LRDs may provide crucial insights into the formation pathways and duty cycles of LRDs.
@@
The situation is further complicated by the lack of a clear definition and robust selection methodology for LRDs in the literature. A commonly adopted definition identifies LRDs as sources with compact morphologies at long wavelengths and V-shape SEDs. Some studies further require the presence of broad emission lines as part of the definition. In practice, many sources are classified as LRDs based on simple color-cut or slope-cut criteria (e.g., 2024arXiv240610341A; 2024arXiv240403576K; 2024ApJ...968...38K; 2025ApJ...978...92L). These photometric colors and fitted slopes may be contaminated by strong emission lines, such as H, H, and [O iii] line (e.g., 2025ApJ...979..138H). This leads to misclassifications or the inclusion of unrelated sources in LRD samples. Therefore, it is crucial to remove the contamination of the line emission when constructing a LRD sample.
@@
In this work, we present a systematic investigation of narrow-line LRDs based on a clean sample selected using emission-line-free photometry. The contamination from strong emission lines is removed using spectroscopic information from NIRSpec/PRISM observations. The narrow H line is identified with high- or medium-resolution NIRSpec grating spectra. We compare the properties of narrow-line LRDs with those of broad-line LRDs and normal star-forming galaxies, aiming to understand the diversity within the LRD population and the possible links between them. The paper is organized as follows. In Section II, we describe our data and data reduction. In Section III, we introduce our selection and characterization of the narrow-line LRD sample. Two possible interpretations of the narrow-line LRDs are proposed in Section IV. In Section V, we discuss the comparison of the sample selection methods and the implication of narrow-line LRDs. Conclusions are given in Section VI. Throughout this paper, we use a cosmology with , , and (2020A&A...641A...6P). By default, measurement uncertainties are quoted at a 1 confidence level, and upper limits are quoted at a 90% confidence level.
II Data Description and Reduction
II.1 Imaging Data and Catalog
@@
We collect publicly available JWST/NIRCam data in A2744, CEERS, GOODS-S, GOODS-N, COSMOS, and UDS fields. The detailed reduction steps can be found in 2025ApJ...985..119Z. Briefly, we reduce the raw data using the combination of the JWST Calibration Pipeline (v.1.12.5; 2023zndo..10022973B), the CEERS NIRCam imaging reduction (2023ApJ...946L..12B),111https://github.com/ceers/ceers-nircam. and our own custom codes. All NIRCam images are aligned to the HST images in each field. We use the public HST image from the CANDELS and 3D-HST surveys (2011ApJS..197...35G; 2011ApJS..197...36K; 2012ApJS..200...13B; 2016ApJS..225...27M). The final mosaics in all fields have pixel scales of .
@@
The empirical PSF models of each field are constructed from the mosaic images with PSFEx (2011ASPC..442..435B), following the method in 2024ApJ...962..139Z. We then use these PSFs to derive kernels to match all PSFs to the F444W PSF using PyPHER (2016ascl.soft09022B). These kernels are used to convolve all images to match the F444W image. Then we perform photometry using SExtractor (1996A&AS..117..393B) on the PSF-matched images to construct source catalog. We use the dual image mode and the detection image is an image combining the PSF-matched F277W, F356W and F444W images. We adopt a two-step ‘cold + hot’ mode to achieve an optimal source detection (2012MNRAS.422..449B). Photometry is measured in a small Kron aperture (=1.1, =1.6), and an aperture correction is performed to match the default Kron aperture (=2.5, =3.5) in each band. Following the method in 2023ApJS..269...16R, we measure the photometric errors by placing 100,000 random apertures in the source-free region of the images. These apertures are divided into 100 groups based on their weight values. For each group, a power-law relation between aperture size and flux scatter is fitted and used to derive the sky background noise for each source according to its weight value and aperture size.
@@
For the CEERS field, we adopt the ASTRODEEP catalog (2024A&A...691A.240M) that is based on the images from the CEERS data release (2023ApJ...946L..12B), since this catalog includes more HST bands. Their basic procedure for photometry is similar to ours, with the major differences outlined below. Source detection was carried out using SExtractor on a weighted combination of the F356W and F444W images. The total fluxes were inferred within a preferred aperture, determined via the object’s segmentation map. The aperture correction in the detection band was applied to other bands under the assumption of no color gradient.
II.2 NIRSpec Spectroscopic Data
@@
The NIRSpec spectroscopic data used in this work are from the DAWN JWST Archive (DJA222https://dawn-cph.github.io/dja/.), which includes fully reduced 1D and 2D spectra of all public JWST/NIRSpec data. We use the version 3 publicly released NIRSpec datasets, including 26,077 spectra for 14,909 sources from 43 programs. The source IDs used in this paper are the UIDs of the DJA dataset. These data are reduced using the latest version of msaexp and the JWST Calibration Pipeline (v.1.14.0) with calibration files jwst_1225.pmap from the Calibration Reference Data System. The detailed MSA data reduction procedures using msaexp can be found in deGraaff_2025_RUBIES and 2024Sci...384..890H. The absolute flux calibration accuracy for NIRSpec MSA spectroscopy is generally at a level of 10% to 20% (deGraaff_2025_RUBIES). DJA also offers spectroscopic redshifts obtained through the template fitting algorithms employed in the msaexp pipeline. These redshifts are visually inspected and graded for reliability, with grade 3 indicating a secure measurement. In our analysis, we only include sources with grade 3 redshift, and adopt the grade 3 redshift as their redshift value.
@@
We also notice the availability of two large NIRSpec datasets from the data releases of two JWST Cycle 1 programs: JWST Advanced Deep Extragalactic Survey (JADES; 2023arXiv230602465E; d2024jades) and the CEERS (2023ApJ...946L..12B). These datasets are carefully reduced and generally offer higher data quality compared to the DJA data (e.g., 2025ApJ...980...93J). For uniformity, we adopt only the DJA data throughout this work. We have verified that this choice has a negligible effect on our results. Most of the CEERS spectra are also included in DJA. In the JADES release, 1092 out of 2525 targets are not available in DJA; however, we identified only 4 LRDs (and no narrow-line LRDs) among them, which may be attributed to the target selection strategy of the JADES survey. Overall, the results obtained from the JADES/CEERS and DJA datasets are consistent for common targets.
III Narrow-line LRDs
III.1 Selection of LRDs
@@
Photometric selection of LRDs in the literature primarily relies on their observed colors to identify the characteristic V-shape SEDs (e.g., 2024arXiv240610341A; 2024ApJ...968...38K; 2025ApJ...978...92L). 2024arXiv240403576K further refined this approach by applying cuts to the rest-frame optical and UV slopes. This photometry-only selection may suffer from inaccurate photo- or boosted red colors by strong emission lines (e.g., 2025ApJ...979..138H).
@@
LRDs are then selected using the criteria of and , as done by 2024arXiv240403576K. We present two examples in Figure 1 illustrating sources with a fake and a genuine V-shape continuum, respectively. For comparison, we also show the LRD template constructed from the average SED of LRDs in 2024arXiv240610341A. In Figure 1(a), the red color of the source is primarily due to its strong H and [O iii] emission lines. After subtracting their contributions, the optical slope becomes flat. In contrast, the source shown in Figure 1(b) remains its V-shape after the emission line contribute is subtracted. The distribution of the best-fit UV and optical slopes for our sample is shown in Figure 2(a). In total, there are 179 sources with and .
To select sources with compact morphology, we apply an additional cut on the half-light radius measured by SExtractor in the F444W-band images. We derive a relation between and for point sources, as shown in Figure 2(b). Sources with below 1.5 times the relation (green dashed line in Figure 2b) are selected. A total of 98 LRDs meet this criterion. The redshift-magnitude diagrams (F115W and F444W) of the selected LRDs are shown in Figure 3. These LRDs span a large redshift range of –9 and a broad magnitude range of –28 mag. In the rest-frame UV, their magnitudes are comparable to those of typical star-forming galaxies reported by 2024A&A...684A..75C, whereas in the rest-frame optical, the LRDs appear brighter.
@@
III.2 Selection of Narrow-line LRDs
We aim to search LRDs that do not have a broad H line component. Among the 98 selected LRDs, 32 have high- or medium-resolution grating spectra covering H, and 20 of them have additional high- or medium-resolution grating spectra covering H and O iii. These LRDs with both PRISM and grating spectra are mainly from the RUBIES (deGraaff_2025_RUBIES), UNCOVER (2024ApJ...974...92B), and JADES (2023arXiv230602465E) programs. We fit these lines using lmfit (2021zndo....598352N). The line spread functions (LSFs) provided by JDOX significantly underestimate the actual spectral resolution (e.g., 2024A&A...684A..87D; 2024A&A...691A.145M), since they apply only for uniformly illuminated sources. Therefore, we use msafit to model the LSF, which forward models the light trajectory on the JWST detector for a given source position to measure the spectral resolution (2024A&A...684A..87D). Some sources have multiple high- or medium-resolution spectra. In such cases, we use the spectrum with the highest S/N rather than combining them to avoid ambiguity due to the LSF effect. The results from different spectra of the same sources are consistent.
For O iii, we use a single Gaussian to fit either line separately. We do not apply any constraint (e.g., 2025ApJ...980...93J). For H and H, we perform two types of fitting: